# Life Finds A Way: Emergence of Cooperative Structures in Adaptive Threshold Networks

### v2

@@ 4.5 Binding Range as Ecological Analog
The binding range parameter, which can be though of as the probability or “ease" of forming interactions (akin to biochemical or ecological compatibility), accelerates the dynamics of community structuring. It is interesting this parameter does not radically alter long-term outcomes but modulates the speed and slightly reduces the SCC’s ceiling under antagonistic conditions. In ecological contexts, this could suggest ecosystems with higher interaction potentials (dense signaling, physical proximity, etc) might achieve stable cooperation more rapidly and sustain slightly larger cohesive structures despite competitive pressures. Nevertheless, even in cases with a low binding range, it is “just a matter of time” before cooperative communities (autocatalytic sets, ecosystems (biological, economic, cultural), etc.) will be formed.
@@ 5 Conclusions and Future Work
We have demonstrated that, under minimal assumptions, adaptive threshold networks naturally give rise to communities of support and cooperation, even in the presence of substantial antagonism. The primary insight is that the vast diversity of possible interactions enables complex systems to reliably find paths towards sustained cooperation. The two phase transitions identified in the model suggest that at some point quantity turns into quality. Thus, our results provide support for the conjecture that life inherently "finds a way," tinkering to build cooperative structures despite adverse conditions.
@@
We can assume that evolution will favor a balanced $\rho$ ratio: too much harm is not sustainable. However, only help might hinder adaptability and evolvability  [16]. To explore this hypothesis systematically, we plan to introduce evolutionary feedback explicitly in the next iteration of our model. Edge weights will dynamically evolve, strengthening or weakening interactions probabilistically over time, reflecting real-world processes of mutual adaptation and co-evolution. Preliminary results and theory suggest this enhancement should lead to rich dynamics, potentially revealing other novel phase transitions or critical thresholds that influence community resilience and connection.
@@ Acknowledgements
We thank Andrea Roli, and Jérome Michaud, for their insightful comments and valuable feedback, which greatly improved the manuscript.

### v3

@@ 4.5 Binding Range as Ecological Analog
The binding range parameter, which can be though of as the probability or “ease” of forming interactions (akin to biochemical or ecological compatibility), accelerates the dynamics of community structuring. It is interesting this parameter does not radically alter long-term outcomes but modulates the speed and slightly reduces the SCC’s ceiling under antagonistic conditions. In ecological contexts, this could suggest ecosystems with higher interaction potentials (dense signaling, physical proximity, etc) might achieve stable cooperation more rapidly and sustain slightly larger cohesive structures despite competitive pressures. Nevertheless, even in cases with a low binding range, it is “just a matter of time” before cooperative communities (autocatalytic sets, ecosystems (biological, economic, cultural), etc.) will be formed.
@@ 5 Conclusions and Future Work
We have demonstrated that, under minimal assumptions, adaptive threshold networks naturally give rise to communities of support and cooperation, even in the presence of substantial antagonism. The primary insight is that the vast diversity of possible interactions enables complex systems to reliably find paths towards sustained cooperation. The two phase transitions identified in the model suggest that at some point quantity turns into quality. Thus, our results provide support for the conjecture that life inherently ”finds a way,” tinkering to build cooperative structures despite adverse conditions.
@@
We can assume that evolution will favor a balanced $\rho$ ratio: too much harm is not sustainable. (This perspective is consistent with empirical studies of Streptomyces communities in soil, where dense networks of antagonistic and growth-promoting interactions coexist, exhibiting balanced interaction probabilities, strong reciprocity, and rapid evolutionary turnover rather than static equilibrium [28].) However, only help might hinder adaptability and evolvability  [15]. To explore this hypothesis systematically, we plan to introduce evolutionary feedback explicitly in the next iteration of our model. Edge weights will dynamically evolve, strengthening or weakening interactions probabilistically over time, reflecting real-world processes of mutual adaptation and co-evolution. Preliminary results and theory suggest this enhancement should lead to rich dynamics, potentially revealing other novel phase transitions or critical thresholds that influence community resilience and connection.
@@ Acknowledgements
We thank Andrea Roli and Jérome Michaud for their insightful comments and valuable feedback, which greatly improved the manuscript.